apply the convolution theorem to find the inverse laplace transform of ;

F(s)= 1 /s(s^2+4s+5)
Answer #1

F(s)= 1 /s(s^2+4s+5)

F(s)= 1 /s(s^2+4s+4+1)

F(s)= 1 /s((s+2)^2+1)

F(s)= (1 /s) * (1/((s+2)^2+1)

we know that L (u(t) = 1/s L(e^-2t sint u(t)) = (1/((s+2)^2+1)

thus

inverse laplace transform given by = u(t)*e^-2t sint u(t)

answered by: chima